Robinson Crusoe can bake 10 loaves of bread (B) in one hour or peel 20 potatoes (P). Friday can bake 5 loaves of bread (B) in an hour or peel 30 potatoes (P).


a) What is the opportunity cost of producing 1 unit of bread and 1 unit of potato by Robison Crusoe and Friday?


Answer _____________________________________________________________________


b) Write PPF equation for Robinson Crusoe and Friday


Answer _____________________________________________________________________


c) If perfect competition prevails (Price = marginal cost), determine the price of each unit of bread and potato.


Answer _____________________________________________________________________


d) Determine comparative advantage of Robinson Crusoe and Friday.


1
Expert's answer
2020-09-02T11:38:01-0400

Solution:

a.). What is the opportunity cost of producing 1 unit of bread and 1 unit of potato by Robison Crusoe and Friday?

The opportunity cost of producing 1 unit of bread by Robinson Crusoe"= \frac{20}{10}= 2 \;potatoes"

The opportunity cost of producing 1 unit of bread by Friday "= \frac{30}{5} = 6 \;potatoes" = 30/5

The opportunity cost of producing 1 unit of potato by Robinson Crusoe"= \frac{10}{20} = \frac{1}{2}\;bread"

The opportunity cost of producing 1 unit of potato by Friday "= \frac{5}{30}=\frac{1}{6} \;bread"

 

b.). The PPF equations for Robinson:

"B=10-2p"

"P=20-\frac{1}{2}b"

      The PPF equations for Friday:

"B=10-6p"

"P=30-\frac{1}{6}b"


c.). In perfect competition, the unit price of bread and potato will be:

     Robinson: Bread "=\frac{1}{2}"     Potato "=2"

     

Friday:      Bread "=\frac{1}{6}"      Potato "=6"


d.). Robinson has the comparative advantage of producing breads since it must give up only 2 potatoes to produce a bread while Friday must give up 6 potatoes to produce a bread.

Friday has the comparative advantage to produce potato since it must give up only "\frac{1}{6}" bread to produce a potato while Robinson has to give up "\frac{1}{2}" bread to produce a potato.